Prove that the set Z of all integers with binary operation * defined by a*b=a+b+1



for all a, b belonging to G is an Abelian group.


1
Expert's answer
2021-12-27T18:33:57-0500

Let us prove that the set "\Z" of all integers with binary operation * defined by "a*b=a+b+1"

for all "a, b \in G" is an Abelian group.


Since for any "a, b \in G" we have that "a*b=a+b+1\in \Z," the operation "*" is defined on the set "\Z."

Taking into account that

"(a*b)*c=(a+b+1)*c=(a+b+1)+c+1=a+(b+c+1)+1\\=a+(b*c)+1=a*(b*c)"

for all "a, b \in G", we conclude that the operation "*" is associative on the set "\Z."

Since for "e=-1" we get that

"-1*a=-1+a+1=a" and "a*(-1)=a-1+1=a"

for each "a\in G," we conclude that "e=-1" is the identity of the semigroup "(\Z,*)."

Taking into account that for any "a\in \Z" and "a^{-1}=-a-2\in \Z" we have that

"a*a^{-1}=a-a-2+1=-1=e" and "a^{-1}*a=-a-2+a+1=-1=e,"

we conclude that "a^{-1}" is the inverse of "a."

Therefore, "(\Z,*)" is a group.

Since "a*b=a+b+1=b+a+1=b*a" for any "a,b\in G," we conclude that "(\Z,*)" is an Abelian group.